40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 16 16:00:01 2019, gautam, Update, CDS, c1iscaux Supermicro setup IMG_7769.JPG
    Reply  Wed Jul 17 21:22:41 2019, gautam, Update, CDS, CM board Latch Enable subtlety 
       Reply  Sun Jul 21 12:55:38 2019, gautam, Update, CDS, CM board Latch Enable test script LatchLogic.pdfLatchLogicTest.png
          Reply  Fri Aug 9 08:59:04 2019, gautam, Update, CDS, Prep for install of c1iscaux 
             Reply  Sun Aug 11 11:47:42 2019, gautam, Update, CDS, Bench test of c1iscaux 
                Reply  Mon Aug 12 17:36:04 2019, gautam, Update, CDS, More bench test of c1iscaux iscauxCheclist.pdf
                   Reply  Mon Aug 12 21:25:19 2019, Koji, Update, CDS, More bench test of c1iscaux 
                   Reply  Tue Aug 13 08:07:09 2019, gautam, Update, CDS, P1--->P2 
                      Reply  Tue Aug 13 14:36:17 2019, gautam, Update, CDS, P1--->P2 
                   Reply  Fri Aug 16 16:40:04 2019, gautam, Update, CDS, 1Y3 work newLook1Y3.JPGIMG_7803.JPGc1lsc_crashed.png
                      Reply  Sat Aug 17 16:49:23 2019, gautam, Update, CDS, More 1Y3 work Screen_Shot_2019-08-17_at_3.00.57_PM.pngScreen_Shot_2019-08-17_at_3.12.23_PM.pngIMG_7804.JPGScreenshot_from_2019-08-17_17-04-47.png
                         Reply  Mon Aug 19 14:36:21 2019, gautam, Update, CDS, c1iscaux remaining work caseForSmallerFootprint.pdf
                            Reply  Tue Aug 20 19:05:24 2019, Koji, Update, CDS, MC1 (and MC3) troubleshoot Screenshot_from_2019-08-20_17-26-01.pngScreenshot_from_2019-08-20_17-43-03.png
                               Reply  Thu Aug 22 12:54:06 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 7x
                                  Reply  Thu Aug 22 20:56:51 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 
                            Reply  Fri Aug 23 18:46:17 2019, Jon, Update, CDS, c1iscaux remaining work 
                               Reply  Sun Aug 25 14:18:08 2019, gautam, Update, CDS, c1iscaux remaining work Screen_Shot_2019-08-25_at_10.38.37_PM.png
    Reply  Thu Jul 18 10:46:04 2019, gautam, Update, CDS, Database files made 
       Reply  Thu Jul 18 19:58:56 2019, gautam, Update, CDS, Work on Acromag chassis IMG_7771.JPGIMG_7770.JPGIMG_7772.JPG
          Reply  Fri Jul 19 19:44:03 2019, gautam, Update, CDS, Database file test Whitening.png
Message ID: 14781     Entry time: Fri Jul 19 19:44:03 2019     In reply to: 14773
Author: gautam 
Type: Update 
Category: CDS 
Subject: Database file test 

Summary:

The database files for C1ISCAUX seem to work file - the exception being the mbbo channels for the CM board.

Details:

This was just a software test - the actual functionality of the channels will have to be tested once the Acromag crate has been installed in the rack. One change I had to make on the MEDM screen for the LSC PD whitening gains was to get rid of the "NMS" suffix on the EPICS channel names for whitening gain sliders/drop-down-menus. I suspect this has to do with the EPICS version we are using, 7.0.1. Furthermore, AS165 and POP55 no longer exist - I hold off removing them from the MEDM screen for the moment.

Next steps:

From the software point of view, the major steps are:

  1. Fix the mbbo channel notation in the database files
  2. Write and test the latch enabling code
  3. Figure out what scripted tests can be done to test the functionality of the new Acromag box.

I am stopping the EPICS server on the new machine and restarting the old VME crate over the weekend.

Attachment 1: Whitening.png  29 kB  Uploaded Fri Jul 19 20:49:10 2019  | Hide | Hide all
Whitening.png
ELOG V3.1.3-